# Probability-Aware Parking Selection

### v1

@@ V-B Strategy comparison
To validate the relative merits of probability-aware parking selection with stochastic observations, we conduct comparisons against traditional (baseline) methods. This work will use simulations of parking trials.
@@ V-B1 Sites
  • (Dense) Area A.2: Similar to Area A.1., Pike Place Market is the destination. The three lots, which comprise 98 spots, are those nearest the destination along Pike Place (Lot 1, closest), 2nd Avenue (Lot 2, next closest), and Alaskan Way (Lot 3, furthest). The origin is Kerry Park, a residential neighborhood chosen for its centrality within the city. Google Maps estimates an uncongested ‘time to drive’ of ten minutes and public transit times of 18-22 minutes (including walk time), depending on the time of day.
  • (Sparse) Area B.2: The Botanical Garden is the destination. The lots, which comprise 69 spots, include one on 54th St. (Lot 1) and the two nearest the Garden on Market St. (Lots 2 and 3, respectively). Again, Lot 1 is the nearest to the destination and Lot 3 is furthest. The origin is Salmon Bay Park, selected to provide geographic heterogeneity as a residential hub in outer Seattle. Google Maps estimates an uncongested ‘time to drive’ of six minutes and public transit times of 21-23 minutes (including walk time), depending on the time of day.
@@ V-B2 Scenarios
To capture temporal heterogeneity, we simulate multiple days and departure times. The weekday selected is Thursday, January 30th, 2025, and the weekend day selected is Saturday, February 1st, 2025. The selection of this weekend date in Area A.2 is intentional to consider the role of event interference: an annual festival brought crowds to the area for music, art, and shopping.
@@ V-B3 Policies
Greedy patient and impatient policies serve as baselines. These are derived from the recommendations of popular navigation apps and how users may behave in response, either by waiting or seeking additional nearby parking [arora2019hard].
@@
  • Baseline-Impatient is an impatient policy that first attempts to park at the lot closest to the destination. If this fails, it drives to the unvisited lot closest from its current position to continue its search. If all lots have been visited, the search cycle resets.
@@ V-B4 Results
For each simulation we compute the ‘time to arrive’ – the simulated time to successfully park and then walk to the destination. This is the per-episode cumulative reward. The results from these experiments are shown in terms of time savings and loss in Table I. The means and standard deviations (in minutes) of trip durations are presented for each setting. Gains are computed as percent improvements over the baseline policies’ mean times. Policies in bold indicate a top performance in either the 10% or the 50% adoption rate category. To better understand the causes of errors and their impacts on the system, we also compare probability-aware policies against oracle baselines that have access to the current true parking probabilities at each lot, rather than those generated from connected user observations. These quantify the loss due to the suboptimal probability awareness.
@@
The findings are notable in comparison to each location’s direct drive time (without parking) and public transit times. Even for the best probability-aware policy in each scenario studied, the average ‘time to arrive’ accounting for the parking search and post-parking walk is 67% to 123% higher than the ‘time to drive’ directly to the destinations. See Table II. This shows how travel time projections that do not account for parking can substantially underestimate the true value.
This difference may be relevant for users deciding which mode to take. In the case of the route from Kerry Park to Pike Place Market (A.2), the travel time estimates for using public transit are similar to the mean travel times of the best probability-aware strategies and substantially shorter than the baseline strategies; see Table III. In contrast, the original ‘time to drive’ projection was 50% shorter than the average public transit time estimate. This could mean that users seeking short travel times may mistakenly disregard public transit because they believe it requires double the time of driving a personal vehicle, when the opposite may be true (or the options may at least be similar). Note that uncongested drive time estimates were used for conservatism – we expect increased traffic would tend to add more time to the parking strategies than the direct drive time, since they generally comprise longer distances.
TABLE II: Comparison of ‘Time to Arrive’ and ‘Time to Drive’ Estimates.
% over Time to Drive (↓)
Destination Day Policy 10% Adoption 50% Adoption
Area A.2 (Pike Place Market) Weekday BL-Patient 346% 352%
BL-Impat. 119% 108%
Best PA 90% 91%
Weekend BL-Patient 456% 462%
BL-Impat. 173% 153%
Best PA 119% 91%
Area B.2 (Botanical Garden) Weekday BL-Patient 72% 70%
BL-Impat. 80% 75%
Best PA 67% 68%
Weekend BL-Patient 143% 133%
BL-Impat. 130% 133%
Best PA 123% 93%
Each policy’s ‘time to arrive’ is the mean recorded in Table I. ‘Best PA’ refers to the probability-aware policy with the lowest mean time for a given scenario. Each destination’s ‘time to drive’ is as described in Section V-B.1. Lower is better.
TABLE III: Comparison of ‘Time to Arrive’ Relative to Public Transit Travel Time Estimate for Area A.2 (Pike Place Market).
% over Public Transit Travel Time
Day Policy 10% Adoption 50% Adoption
Weekday BL-Patient 123% 126%
BL-Impat. 10% 4%
Best PA -5% -4%
Weekend BL-Patient 178% 181%
BL-Impat. 37% 27%
Best PA 10% -4%
Each policy’s ‘time to arrive’ estimate is its mean from Table I. ‘Best PA’ is defined as in Table II. The public transit travel time estimate used is 20 minutes, the midpoint of the range described in the Section V-B.1 site description for Area A.2. For reference, this site’s ‘time to drive’ estimate (10 minutes) is 50% lower than the estimated public transit time.
@@ V-B5 Error impact
Interestingly, performance is comparable across the adoption rates. This suggests limited operational impact from errors incurred via stochastic observation. The probability-aware policies’ performance relative to their oracles also supports this conclusion, although those at the 50% adoption rate perform slightly better relative to their oracles than those at the 10% rate. One explanation for the system’s resilience to errors may be the following trade-off: probability estimates (and by extension, errors) may matter more in competitive settings, but competitive parking environments also have higher traffic, which produces more frequent observations of a lot’s parking availability. The results across the 1-, 2-, and 3-step lookahead probability-aware policies suggest that deeper lookaheads – which incur computational complexity exponential in the depth of the search – may not be necessary.
@@ VI Future work
Future work could find analytical expressions for alternative cases or explore uncertainty in drive and walk time estimates. Large-scale experiments assessing the advantages and disadvantages of the strategies outlined above would be welcome, as would identification of parking clusters $\mathcal{C}$ and investigation of additional probability-aware strategies. Related work could be done to estimate the reduction in carbon emissions and time saved via mode shift if ‘time to drive’ estimates are corrected to ‘time to arrive’ estimates that take parking availability and post-parking walking into account. These could also integrate driver preference models that consider factors beyond travel time alone.
[Uncaptioned image] Cameron Hickert earned a B.S. in physics from the University of Denver and an M.E. in computational science and engineering from Harvard University. He is currently a Ph.D. student in Social and Engineering Systems at MIT. His research interests include safe autonomy, reinforcement learning, and distributed cyber-physical systems.
[Uncaptioned image] Sirui Li received her B.S. degree with majors in computer science and mathematics from Washington University in St. Louis, MO, USA, in 2019. She is currently working toward her Ph.D. degree in Social and Engineering Systems at MIT, Cambridge, MA, USA. Her research interests include areas of machine learning for combinatorial optimization and control analysis for transportation systems.
[Uncaptioned image] Zhengbing He (M’17-SM’20) received the Bachelor of Arts degree in English language and literature from Dalian University of Foreign Languages, China, in 2006, and the Ph.D. degree in systems engineering from Tianjin University, China, in 2011. He was a Post-Doctoral Researcher and an Assistant Professor with Beijing Jiaotong University, China. From 2018 to 2022, he was a Full Professor with Beijing University of Technology, China. Presently, he is a Research Scientist at MIT. His research stands at the intersection of transportation, systems engineering, and artificial intelligence. He has published more than 150 academic papers, with total citations exceeding 6,000. He is the Editor-in-Chief of Journal of Transportation Engineering and Information (Chinese). Meanwhile, he serves as a Senior Editor for IEEE TRANSACTIONS ON INTELLIGENT TRANSPORTATION SYSTEMS, an Associate Editor for IEEE TRANSACTIONS ON INTELLIGENT VEHICLES, a Handling Editor for Transportation Research Record, and an Editorial Advisory Board Member for Transportation Research Part C. His webpage is https://www.GoTrafficGo.com.
[Uncaptioned image] Cathy Wu is the Thomas D. and Virginia W Cabot Career Development Associate Professor of Civil and Environmental Engineering at MIT and holds additional appointments in LIDS and IDSS. She holds a Ph.D. from UC Berkeley and B.S. and M.Eng. from MIT, all in EECS, and completed a Postdoc at Microsoft Research. Her interests are at the intersection of machine learning, control, and mobility. Her recent research focuses on how learning-enabled methods can better cope with the complexity, diversity, and scale of control and operations in mobility systems. She is interested in developing principled computational tools to enable reliable decision-making in sociotechnical systems.

### v2

@@ V-B Strategy comparison
To validate the relative merits of probability-aware parking selection with stochastic observations, we conduct comparisons against traditional (baseline) methods. This uses simulations of parking trials.
@@ V-B1 Sites
  • (Dense) Area A.2: Similar to Area A.1., Pike Place Market is the destination. The three lots, which comprise 98 spots, are those nearest the destination along Pike Place (Lot 1, closest), 2nd Avenue (Lot 2, next closest), and Alaskan Way (Lot 3, furthest). The origin is Kerry Park, a residential neighborhood chosen for its centrality within the city. Google Maps estimates an uncongested time-to-drive of ten minutes and public transit total travel times (time-to-arrive) of 18-22 minutes, depending on the time of day.
  • (Sparse) Area B.2: The Botanical Garden is the destination. The lots, which comprise 69 spots, include one on 54th St. (Lot 1) and the two nearest the Garden on Market St. (Lots 2 and 3, respectively). Again, Lot 1 is the nearest to the destination and Lot 3 is furthest. The origin is Salmon Bay Park, selected to provide geographic heterogeneity as a residential hub in outer Seattle. Google Maps estimates an uncongested time-to-drive of six minutes and public transit times of 21-23 minutes (again including walk time).
@@ V-B2 Scenarios
To capture temporal heterogeneity, we simulate multiple days and departure times. The weekday selected is Thurs., Jan. 30th, 2025, and the weekend day selected is Sat., Feb. 1st, 2025. The selection of this weekend date for Area A.2 is intentional to consider event interference: a festival brought crowds to the area for music, art, and shopping.
@@ V-B3 Policies
Greedy patient and impatient policies serve as baselines. These are derived from the recommendations of popular navigation apps and how users may behave in response, either by waiting or seeking additional nearby parking [2].
@@
  • Baseline-Impatient is an impatient policy that first attempts to park at the lot closest to the destination. If this fails, it drives to the unvisited lot closest to its current position to continue its search. If all lots have been visited, the search cycle resets.
@@ V-B4 Results
For each simulation we compute the time-to-arrive. This is the per-episode cumulative reward, as described in Section III-A. The results from these experiments are shown in terms of time savings and loss in Table I. The means and standard deviations (in minutes) of trip durations are presented for each setting. Gains are computed as percent improvements over the baseline policies’ mean times. Policies in bold indicate a top performance in either the 10% or the 50% adoption rate category. To better understand the causes of errors and their impacts on the system, we also compare probability-aware policies against oracle baselines that have access to the current true parking probabilities at each lot, rather than those generated from connected user observations. These quantify the effect of suboptimal probability awareness.
@@
The findings are notable in comparison to each location’s time-to-drive. See Table II. Even for the best probability-aware policies across all scenarios studied, the average time-to-arrive is 67% to 123% higher than the time-to-drive directly to the destinations. This discrepancy is up to 462% and 173% for the patient and impatient baselines, respectively. This shows how travel time projections that do not account for parking can substantially underestimate the true value.
This difference may be relevant for users deciding which mode to take. In the case of the route from Kerry Park to Pike Place Market (A.2), the public transit time-to-arrive estimates are substantially shorter than the mean time-to-arrive values of the baseline driving strategies and similar to the values from the best probability-aware driving strategies; see Table III. In contrast, the time-to-drive projection is 50% shorter than the average public transit time-to-arrive. This could mean that users may mistakenly disregard public transit because they believe it requires double the time of driving a personal vehicle, when the opposite may be true (or the options may at least be similar). Note that uncongested drive time estimates were used for conservatism – we expect increased traffic would tend to add more time to the parking strategies than the direct drive time, since they generally comprise longer distances.
TABLE II: Time-to-Arrive vs. Time-to-Drive: Abs. & Pct. Differences (↓)
10% Adoption 50% Adoption
Dest. Day Policy Minutes % Minutes %
[…]
Values represent the absolute gap and percentage increase of time-to-arrive over time-to-drive. Each policy’s time-to-arrive is the mean recorded in Table I. ‘Best PA’ refers to the probability-aware policy with the lowest mean time for a given scenario. Each destination’s time-to-drive is as described in Section V-B.1. Lower is better. Darker shading indicates more severe time-to-drive underestimates relative to time-to-arrive.
TABLE III: Driving Time-to-Arrive vs. Public Transit Time-to-Arrive for Area A.2 (Pike Place Market): Abs. & Pct. Differences
10% Adoption 50% Adoption
Day Policy Minutes % Minutes %
[…]
Values represent the difference in time-to-arrive for driving relative to public transit. Each driving policy’s time-to-arrive is defined as in Table II. The public transit time-to-arrive estimate used is 20 minutes, the midpoint of the range described in the Section V-B.1 site description for Area A.2. For reference, this site’s time-to-drive estimate is 10 minutes shorter (-50%) than the estimated public transit time-to-arrive. Darker shading indicates larger deviations. Purple is used since larger values are not necessarily worse – more accurate total travel time estimates enable fairer cross-mode comparison.
@@ V-B5 Error impact
Interestingly, performance is comparable across the two adoption rates. This suggests limited operational impact from errors incurred via stochastic observation. The probability-aware policies’ performance relative to their oracles also supports this conclusion, although those at the 50% adoption rate perform slightly better relative to their oracles than those at the 10% rate. One explanation for the system’s resilience to errors may be the following trade-off: probability estimates (and by extension, errors) may matter more in competitive settings, but competitive parking environments also have higher traffic, which produces more frequent observations of a lot’s parking availability. The results across the 1-, 2-, and 3-step lookahead probability-aware policies suggest that deeper lookaheads – which incur computational complexity exponential in the depth of the search – may not be necessary.
@@ VI Future work
Future work could find analytical expressions for alternative cases or explore uncertainty in drive and walk time estimates. Large-scale experiments assessing the advantages and disadvantages of the strategies outlined above would be welcome, as would identification of parking clusters $\mathcal{C}$ and investigation of additional probability-aware strategies, perhaps with reinforcement learning. Related work could be done to estimate the reduction in carbon emissions and time saved via mode shift if time-to-drive estimates are corrected to time-to-arrive estimates. These could also integrate driver preference models that consider factors beyond travel time.